OpenStudy (anonymous):

log8x-log(1+sqrt(x)) = 2

OpenStudy (anonymous):

start with \[\log(\frac{8x}{1+\sqrt{2}})=2\] \[\frac{8x}{1+\sqrt{2}}=100\] \[8x=100+100\sqrt{2}\] \[x=\frac{100+100\sqrt{2}}{8}\]

OpenStudy (anonymous):

assuming this is log base ten

OpenStudy (anonymous):

where did you get the square root of 2 from?

hero (hero):

He multiplied both sides by \[1+\sqrt{2}\]

OpenStudy (anonymous):

thanks for the effort

hero (hero):

On the right side you distribute the 100

hero (hero):

x is approximately 30.18

OpenStudy (anonymous):

Umm...okay. So what you really do. Combine the logs using the property that: \[\log(a)-\log(b)=\log(\frac{a}{b})\] \[\log(\frac{8x}{1+\sqrt{x}})=2\] Raise all of it to the 10^(whatever) Giving: \[\frac{8x}{1+\sqrt{x}}=10^2\] On the left side multiply top and bottom by 1-sqrt(x) giving: \[\frac{8x(1-\sqrt{x})}{(1+\sqrt{x})(1-\sqrt{x})} \rightarrow \frac{8x-8x^{\frac{3}{2}}}{1-x}\] Multiply across: \[8x-8x^{\frac{3}{2}}=100(1-x) \rightarrow 8x-8x^{\frac{3}{2}}=100-100x\] \[8x^{\frac{3}{2}}-108x+100=0\] Using rational roots theorem you can deduce one of the solutions is x=1. Then factor it by long division. http://www.wolframalpha.com/input/?i=solve+8x^(3%2F2)-108x%2B100%3D0

OpenStudy (anonymous):

The real solution is 180 something. The x=1 comes from the x intercept, the other point is the real solution.